Kenny is 14 years older than Chad. Next year Kenny will be 3 times as old as Chad. How old is Kenny?
Could you please give me all the steps to finding this answer.

Answer by venugopalramana(3286)

Kenny is 14 years older than Chad
LET CHAD BE X YEARS...SO KENNY IS X+14 YEARS
. Next year Kenny will be 3 times as old as Chad.
NEXT YEAR CHAD WILL BE X+1 YEARS AND KENNY WILL BE X+14+1=X+15 YRS.
SO X+15=3(X+1)=3X+3
3X-X=15-3=12
2X=12
X=6
SO CHAD IS 6 YEARS AND
How old is Kenny?
KENNY IS 6+14=20 YEARS NOW.